OpenStudy (anonymous):

SOS!! In your own words explain the difference between the following two expressions: (-1)^2/4 and (-1)^1/2

OpenStudy (anonymous):

Negative one to the power of 2/4. Negative one to the power of 1/2. They both equal -1.5. I need the difference!

OpenStudy (owlcoffee):

if we have: \[(-1)^{\frac{ 2 }{ 4 }}\] \[(-1)^{\frac{ 1 }{ 2 }}\] Now there is a property of the exponents that allows me to convert the fractionary ecponents into one number, I will now prove it, but I will show it to you: \[a ^{\frac{ m }{ n }}=\sqrt[n]{a ^{m}}\] "a" is any number that is not zero, "m" and "n" are any real number but "n" cannot be zero. So let's apply it to both cases and see what we get: \[(-1)^{\frac{ 2 }{ 4 }}=\sqrt[4]{(-1)^{2}}\] \[(-1)^{\frac{ 1 }{ 2 }}=\sqrt{(-1)}\] This is the most important part, so we know that any negative number squared becomes positive, and the equare root of -1 is a complex number. so we can write: \[(-1)^{\frac{ 2 }{ 4 }}=\sqrt[4]{(-1)^{2}} = \sqrt[4]{1}=1\] \[(-1)^{\frac{ 1 }{ 2 }}=\sqrt{(-1)} = i\] so we can say that the difference between those two cases is that one gives us a real number and the other an imaginary number.

OpenStudy (anonymous):

Wow thank you!!